OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 Any Real Number No Solution 0 3

OpenStudy (anonymous):

2x+2=2x+5 2x-2x=5-2 0=3 I guess it has no solution :)

OpenStudy (anonymous):

2x gets cancelled and 2=5 is left.(which is absurd) So no solution as said by @xoya .

OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (anonymous):

7x-5x= 10+1+7 2x=18 x=9

OpenStudy (anonymous):

Yes :)